OpenStudy (anonymous):

factors of 45 (smallest to largest)

OpenStudy (anonymous):

1,3,5,9,15,45,

OpenStudy (anonymous):

1*45=45 3*15=45 5*9=45

OpenStudy (anonymous):

what numbers can multiply each other to equal 45? use rainbow method (or something like that.)